# Presenting a STEM Ways of Thinking Framework for Engineering Design-based Physics Problems

### v1

@@ 3.4 The Five WoT4EDP Elements
In our context, mathematics serves as an indispensable tool for solving engineering design-based problems. This perspective is further reinforced by Tiles’ assertion that mathematics “is not just a language which can be translated into and out of, rather it plays an essential role in the articulation of scientific concepts” [53, p.3].
@@
The National Research Council’s How People Learn: Brain, Mind, Experience, and School: Expanded Edition (2000) [56] serves as a foundational document that underscores the significance of metacognition in student learning. The report argues that adopting a metacognitive approach to instruction “can help students learn to take control of their own learning by defining learning goals and monitoring their progress in achieving them” [56, p.18]. This perspective is built on the understanding that metacognitive thinking encourages students to engage in an “internal conversation” [56, p.18] that fosters self-regulation and self-directed learning. Furthermore, it posits that metacognition can empower students to “teach themselves” [56, p.50]. Consequently, the report advocates for the integration of metacognitive skills into various subject areas within the curriculum.
@@
In the context of this report, the lead author ‘reflected’ on the definition of metacognition. Initially, he perceived it simply as “reflecting, or thinking about thinking”. However, he realized this may be a naive view considering Crawford and Capps’ assertion that “Reflection and metacognition are both thinking processes. Yet, reflection and metacognition differ in some ways” [57, p.16]. Our aim in this article is not to engage in a debate regarding the precise definition of metacognition. Rather, it is perhaps prudent to align with Livingston’s assertion that “In actuality, defining metacognition is not that simple” [76, p.2]. Given the complexities surrounding the term, we find it beneficial to draw on the fine example provided by Kohen and Kramarski, who note that “presenting data in a graph is a cognitive function, whereas reflecting on the answer and realizing that the graph fits the givens are part of the metacognitive process” [57, p.280]. For the records, we consider metacognition as a “deliberate, planful, and goal-oriented mental process, using higher level thinking skills applied to one’s thoughts and experiences” [57, p.17].
@@
A compelling reason for incorporating metacognitive reflection as a critical way of thinking within our framework is best articulated in Hacker and Dunlosky’s thoughtful essay, Not All Metacognition Is Created Equal. They emphasize that students “should not be instructed to merely use ‘metacognition’ but instead should be directed toward specific, higher-level analyses in which they explain how they are solving the problem” [58, p.79]. This insistence on directing students toward higher-level analysis highlights the need for metacognitive reflection to be not only present, but also be purposeful in guiding students toward more profound understanding and effective problem-solving strategies.
@@
A similar perspective is echoed in the NRC’s 2012 report, which highlights how metacognitive approaches in STEM education are embedded in problem-based learning and reflective activities designed to promote critical thinking. While instructors may assume students possess these skills or view them as too advanced for introductory courses, research shows that explicitly teaching metacognitive skills alongside science content enhances student responsibility, problem-solving abilities, and study habits [39, p.154].
@@
Another recent study by Ulu and Yerdelen-Damar [80, p.20, 21] employs a metacognition framework inspired by Brown, Schraw, and Dennison [81, 82]. Metacognition refers to understanding and regulating one’s own thinking and is divided into two components: (i) Knowledge of cognition—declarative information about one’s cognitive processes—and (ii) Regulation of cognition—the ability to plan, monitor, control, and evaluate one’s thinking. Knowledge of cognition consists of three levels: declarative knowledge (facts and strategies), procedural knowledge (how to apply strategies), and conditional knowledge (when and why to apply strategies). Regulation of cognition involves five key skills: planning, managing information, monitoring comprehension, debugging, and evaluating. Although we did not adopt this framework in our earlier studies, it holds promise for future integration into our WoT4EDP framework. Its structured approach to analyzing metacognitive thinking aligns with the reflective practices we aim to foster in design-based activities.
@@
Integrating computational practices within STEM classrooms equips learners with an authentic understanding of scientific processes and better prepares them for careers in an increasingly computational world. As computational thinking (CT) becomes essential across STEM disciplines, the demand for computation-integrated physics courses continues to grow, underscoring the notion that computation is synonymous with engaging in STEM. The urgency and significance of incorporating computational thinking (CT) based activities into physics education is powerfully highlighted by Hamerski et al. [83, p.23], who raise a “call to action” in their case study of high school students involved in a computation-integrated physics course. They urge researchers to recognize and respond to the pressing need for integrating computation into educational practices.
@@
Wing’s influential definition of CT as “solving problems, designing systems, and understanding human behavior by drawing on concepts fundamental to computer science” [85, p.33] demonstrates that the concept extends far beyond computer science alone. This definition also establishes the relevance of CT to STEM fields, particularly in the context of engineering design (ED)-based problems, which focus on developing human-centered, science-based solutions for real-life challenges [28]. In our teaching context, we find Berland and Wilensky’s more approachable interpretation—“the ability to think with the computer-as-tool”—to be both compelling and simple [86, p.630].
While educators may choose various computational thinking (CT) frameworks [61, p.151, 152] that best suit their contexts, we have opted for Shute et al.’s framework due to its relevance to our instructional goals. Our students work extensively with Python in the Jupyter Notebook [30] environment via Google Colab [31], which makes Shute et al.’s definition of CT particularly relevant, simple, and practical: “CT as the conceptual foundation required to solve problems effectively and efficiently (i.e., algorithmically, with or without the assistance of computers), with solutions that are reusable in different contexts” [61, p.153]. This CT framework comprises six key facets: (i) Decomposition—breaking a complex system or problem into manageable, functional sub-problems; (ii) Abstraction—extracting key elements, including data collection, pattern recognition, and modeling; (iii) Algorithms—developing logical, step-by-step instructions for solving problems, often represented through flowcharts and implemented in code (e.g., Python); (iv) Debugging—identifying and fixing errors in code or algorithms; (v) Iteration—refining the algorithm or code through repeated improvements; and (vi) Generalization—applying algorithms, code, or CT skills across different domains, enabling efficient problem-solving [61, p.153].
Importantly, Shute et al. emphasize that this framing of CT underpins and is evidenced through a variety of subject areas, “including Mathematics, Science, and even English Language Arts”. This perspective reinforces how various ways of thinking—mathematical, scientific, and computational—are closely linked, facilitating a holistic approach to problem-solving in our educational practices [61, p.146, 151].
@@ 3.5 Pedagogical Goals
In this subsection, we outline the pedagogical goals that this framework can effectively foster. While our goals are rooted in our combined experience as educators and researchers, they are also informed by various reform documents. These goals are neither rigid nor exhaustive; rather, we see them as adaptable to the specific needs of educators and researchers. We recognize that these terms may be interpreted differently across contexts, but we view this diversity as a strength rather than a limitation. Following Gee, we conceptualize these terms as “thinking devices” [88, p.6], offering flexibility in their interpretation and application.
Though learning physics is not solely about applications [89, 90, 91, 92], the Phys21: Preparing Physics Students for 21st Century Careers (2016) report underscores the importance of connecting physics to real-world applications [93, p.26]. This integration, which embodies Integrative Thinking, goes beyond connecting sub-fields within physics, extending to interdisciplinary links among physics, math, biology, and beyond, as well as the various ways of thinking outlined in our framework. The report also highlights concerns that current curricula are lacking in key areas such as contextual understanding of core concepts, real-world problem-solving, collaboration, communication, and technology skills, along with lifelong learning and innovation [93, p.11]. We believe that a deliberate focus on WoT frameworks—not just our own—can help address these concerns.
Problem Solving is a most natural goal for any science educator, yet it is almost taken for granted. The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum emphasize that physics is an approach to problem-solving that relies on direct observation and hands-on experimentation. To succeed in this field, one must integrate a wide range of knowledge and skills, including mathematical, computational, experimental, and practical abilities, while also cultivating a “thinking like a physicist” mindset [50, p.iii]. A strong grasp of scientific principles is important for students to effectively and efficiently solve problems. As the NGSS (2013) rightly notes, “higher-level reasoning and problem-solving practices require a reasonable depth of familiarity with the content of a given scientific topic if students are to engage in them meaningfully” [38, p.15].
Regarding Learning Innovation, we are largely guided by English’s work, and we will return to this in more detail in Section (iv). While innovation in this context refers to students’ learning, we are tempted to extend it as a goal for educators too. We would approach the term ‘innovation’ rather cautiously, and the SPIN-UP Report aptly captures our tempered view: true ‘Eureka’ moments are rare in science, and even rarer in education. Progress in teaching and learning is incremental and ongoing, with innovations needing to be adapted to suit local needs [94, p.49, 67].
@@
“Philosophical inquiry as pedagogy must be integrated into the curriculum as well as integrate the curriculum.” [95, p.1057].
@@ 3.7 Implications of our Framework
“It is important that developers of quality frameworks not fall into the trap of substituting intuition and guesswork for evidence-based validation processes.” [105, p.361]
@@ 4 WoT Frameworks - A Comparative Study
“Each world whilst it is attended to is real after its own fashion; only the reality lapses with the attention.” [115, p.207].
@@ 4.2 English’s Framework
Systems Thinking and Design-Based Thinking are not discussed here to avoid redundancy. See Tables 4 and 2.
@@
(1) Critical Thinking: Several undergraduate STEM reform documents emphasize critical thinking as essential for developing and refining ideas, whether through explanations, designs, or investigations. Engaging in argumentation and critique foster this thinking and deepen understanding [39, p.125],  [38, p.xv], [73, p.46]. Given this emphasis in reform documents, critical thinking rightly finds a place in English’s framework.
@@
While research indicates that argumentation can enhance conceptual learning [118, p.1304, p.1305][119], Osborne expresses concern about its notable absence in science education. He highlights the need for a deeper understanding of how argumentation facilitates learning and the characteristics of learning environments that foster effective student arguments [51]. In our exploration of SWoT frameworks within scientific argumentation and evidence-based reasoning, we found Slavit et al.’s application of Toulmin’s Claim-Evidence-Reasoning (CER) framework [120, 121, 119] in the context of integrated STEM to be particularly insightful [24, p.135][51, p.464]. An examination of Slavit et al.’s analytic coding framework provides valuable insights with potential applications for future qualitative analyses, particularly in applying CER within integrated STEM contexts.
@@
English’s approach is echoed by Stillman et al., who stress the importance of learning environments that enhance students’ social-critical efficacy. They promote a socio-critical mathematical modeling cycle, allowing students to address problems relevant to their interests and communities while producing solutions that are socially significant [123, p.18]. This perspective aligns with current educational practices that emphasize real-world applications and social relevance in mathematics.
In our context, a catapult design project enables students to collaborate in launching a projectile to a specific range while applying principles of projectile motion. This activity incorporates mathematical modeling, allowing students to make assumptions about factors such as launch angles, launch speed, and drag force. They engage in critical discussions regarding material choices and design constraints, fostering argumentation as they justify their decisions. Before creating an actual model, students can run simulations in Python to measure projectile range, which facilitates reflection and iteration as needed. Furthermore, showcasing their models at a community event not only makes the project fun and educational but also extends the reach of science to the broader community.
@@
In this context, Gardner implores us to get comfortable with the truth, emphasizing that true inquiry transcends mere conversation. She cautions educators to differentiate between facilitating dialogue and facilitating genuine inquiry, noting that quite often “just dialogue continues to pass for philosophical inquiry” [125, p.72]. This distinction has prompted us to reflect on our practices critically.
@@ 5 Summary
To address RG2, we conducted an in-depth comparison between our framework and two others: one by Dalal et al. in the context of Engineering Education Research (EER), and the other by English, focusing on Mathematical and STEM-based problem solving. We outlined the context of each of the studies in sections (i) and (i), and elaborated on each framework in sections (ii) and (iii) respectively. It was insightful and illuminating to draw connections between our framework and the other two, particularly in the comparative analyses shown in Tables 4 and 5. While acknowledging the limitations of our comparative analysis in section 4.3, we also identified a few key implications of our comparative study in section $4.$
@@ 6 Future Work
In the second installment of this two-part series, we will utilize our Ways of Thinking in Engineering Design-based Physics (WoT4EDP) framework as a lens to investigate new data from group discussions and written reports in student-generated engineering design projects. This subsequent study aims to explore how students engage with the elements of design, science, mathematics, and metacognitive reflection throughout their problem-solving approaches. Additionally, we will analyze students’ demonstrations of computational thinking, with a specific focus on Python coding, as they tackle real-world challenges. By applying the WoT4EDP framework, we seek to obtain insights into students’ integrated thinking and Learning Innovation, as emphasized by English. This work will not only expand the application of our framework but also contribute to the ongoing discourse in STEM education regarding effective teaching and learning practices.

### v3

@@ 3.4 The Five WoT4EDP Elements
In our context, mathematics serves as an indispensable tool for solving engineering design-based problems. This perspective is further reinforced by Tiles’ assertion that mathematics “is not just a language which can be translated into and out of, rather it plays an essential role in the articulation of scientific concepts” [59, p.3].
@@
The National Research Council’s How People Learn: Brain, Mind, Experience, and School: Expanded Edition (2000) [62] serves as a foundational document that underscores the significance of metacognition in student learning. The report argues that adopting a metacognitive approach to instruction “can help students learn to take control of their own learning by defining learning goals and monitoring their progress in achieving them” [62, p.18]. This perspective is built on the understanding that metacognitive thinking encourages students to engage in an “internal conversation” [62, p.18] that fosters self-regulation and self-directed learning. Furthermore, it posits that metacognition can empower students to “teach themselves” [62, p.50]. Consequently, the report advocates for the integration of metacognitive skills into various subject areas within the curriculum.
@@
In the context of this report, the lead author ‘reflected’ on the definition of metacognition. Initially, he perceived it simply as “reflecting, or thinking about thinking”. However, he realized this may be a naive view considering Crawford and Capps’ assertion that “Reflection and metacognition are both thinking processes. Yet, reflection and metacognition differ in some ways” [63, p.16]. Our aim in this article is not to engage in a debate regarding the precise definition of metacognition. Rather, it is perhaps prudent to align with Livingston’s assertion that “In actuality, defining metacognition is not that simple” [82, p.2]. Given the complexities surrounding the term, we find it beneficial to draw on the fine example provided by Kohen and Kramarski, who note that “presenting data in a graph is a cognitive function, whereas reflecting on the answer and realizing that the graph fits the givens are part of the metacognitive process” [63, p.280]. For the records, we consider metacognition as a “deliberate, planful, and goal-oriented mental process, using higher level thinking skills applied to one’s thoughts and experiences” [63, p.17].
@@
A compelling reason for incorporating metacognitive reflection as a critical way of thinking within our framework is best articulated in Hacker and Dunlosky’s thoughtful essay, Not All Metacognition Is Created Equal. They emphasize that students “should not be instructed to merely use ‘metacognition’ but instead should be directed toward specific, higher-level analyses in which they explain how they are solving the problem” [64, p.79]. This insistence on directing students toward higher-level analysis highlights the need for metacognitive reflection to be not only present, but also be purposeful in guiding students toward more profound understanding and effective problem-solving strategies.
@@
A similar perspective is echoed in the NRC’s 2012 report, which highlights how metacognitive approaches in STEM education are embedded in problem-based learning and reflective activities designed to promote critical thinking. While instructors may assume students possess these skills or view them as too advanced for introductory courses, research shows that explicitly teaching metacognitive skills alongside science content enhances student responsibility, problem-solving abilities, and study habits [44, p.154].
@@
Another recent study by Ulu and Yerdelen-Damar [86, p.20, 21] employs a metacognition framework inspired by Brown, Schraw, and Dennison [87, 88]. Metacognition refers to understanding and regulating one’s own thinking and is divided into two components: (i) Knowledge of cognition—declarative information about one’s cognitive processes—and (ii) Regulation of cognition—the ability to plan, monitor, control, and evaluate one’s thinking. Knowledge of cognition consists of three levels: declarative knowledge (facts and strategies), procedural knowledge (how to apply strategies), and conditional knowledge (when and why to apply strategies). Regulation of cognition involves five key skills: planning, managing information, monitoring comprehension, debugging, and evaluating. Although we did not adopt this framework in our earlier studies, it holds promise for future integration into our WoT4EDP framework. Its structured approach to analyzing metacognitive thinking aligns with the reflective practices we aim to foster in design-based activities.
@@
Integrating computational practices within STEM classrooms equips learners with an authentic understanding of scientific processes and better prepares them for careers in an increasingly computational world. As computational thinking (CT) becomes essential across STEM disciplines, the demand for computation-integrated physics courses continues to grow, underscoring the notion that computation is synonymous with engaging in STEM. The urgency and significance of incorporating computational thinking (CT) based activities into physics education is powerfully highlighted by Hamerski et al. [89, p.23], who raise a “call to action” in their case study of high school students involved in a computation-integrated physics course. They urge researchers to recognize and respond to the pressing need for integrating computation into educational practices.
@@
Wing’s influential definition of CT as “solving problems, designing systems, and understanding human behavior by drawing on concepts fundamental to computer science” [91, p.33] demonstrates that the concept extends far beyond computer science alone. This definition also establishes the relevance of CT to STEM fields, particularly in the context of engineering design (ED)-based problems, which focus on developing human-centered, science-based solutions for real-life challenges [33]. In our teaching context, we find Berland and Wilensky’s more approachable interpretation—“the ability to think with the computer-as-tool”—to be both compelling and simple [92, p.630].
While educators may choose various computational thinking (CT) frameworks [67, p.151, 152] that best suit their contexts, we have opted for Shute et al.’s framework due to its relevance to our instructional goals. Our students work extensively with Python in the Jupyter Notebook [35] environment via Google Colab [36], which makes Shute et al.’s definition of CT particularly relevant, simple, and practical: “CT as the conceptual foundation required to solve problems effectively and efficiently (i.e., algorithmically, with or without the assistance of computers), with solutions that are reusable in different contexts” [67, p.153]. This CT framework comprises six key facets: (i) Decomposition—breaking a complex system or problem into manageable, functional sub-problems; (ii) Abstraction—extracting key elements, including data collection, pattern recognition, and modeling; (iii) Algorithms—developing logical, step-by-step instructions for solving problems, often represented through flowcharts and implemented in code (e.g., Python); (iv) Debugging—identifying and fixing errors in code or algorithms; (v) Iteration—refining the algorithm or code through repeated improvements; and (vi) Generalization—applying algorithms, code, or CT skills across different domains, enabling efficient problem-solving [67, p.153].
Importantly, Shute et al. emphasize that this framing of CT underpins and is evidenced through a variety of subject areas, “including Mathematics, Science, and even English Language Arts”. This perspective reinforces how various ways of thinking—mathematical, scientific, and computational—are closely linked, facilitating a holistic approach to problem-solving in our educational practices [67, p.146, 151].
@@ 3.5 Pedagogical Goals
In this subsection, we outline the pedagogical goals that this framework can effectively foster. While our goals are rooted in our combined experience as educators and researchers, they are also informed by various reform documents. These goals are neither rigid nor exhaustive; rather, we see them as adaptable to the specific needs of educators and researchers. We recognize that these terms may be interpreted differently across contexts, but we view this diversity as a strength rather than a limitation. Following Gee, we conceptualize these terms as “thinking devices” [94, p.6], offering flexibility in their interpretation and application.
Though learning physics is not solely about applications [95, 96, 97, 98], the Phys21: Preparing Physics Students for 21st Century Careers (2016) report underscores the importance of connecting physics to real-world applications [99, p.26]. This integration, which embodies Integrative Thinking, goes beyond connecting sub-fields within physics, extending to interdisciplinary links among physics, math, biology, and beyond, as well as the various ways of thinking outlined in our framework. The report also highlights concerns that current curricula are lacking in key areas such as contextual understanding of core concepts, real-world problem-solving, collaboration, communication, and technology skills, along with lifelong learning and innovation [99, p.11]. We believe that a deliberate focus on WoT frameworks—not just our own—can help address these concerns.
Problem Solving is a most natural goal for any science educator, yet it is almost taken for granted. The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum emphasize that physics is an approach to problem-solving that relies on direct observation and hands-on experimentation. To succeed in this field, one must integrate a wide range of knowledge and skills, including mathematical, computational, experimental, and practical abilities, while also cultivating a “thinking like a physicist” mindset [56, p.iii]. A strong grasp of scientific principles is important for students to effectively and efficiently solve problems. As the Next Generation Science Standards (NGSS) 2013 report rightly notes, “higher-level reasoning and problem-solving practices require a reasonable depth of familiarity with the content of a given scientific topic if students are to engage in them meaningfully” [43, p.15].
Regarding Learning Innovation, we are largely guided by English’s work, and we will return to this in more detail in Section (iv). While innovation in this context refers to students’ learning, we are tempted to extend it as a goal for educators too. We would approach the term ‘innovation’ rather cautiously, and the SPIN-UP Report aptly captures our tempered view: true ‘Eureka’ moments are rare in science, and even rarer in education. Progress in teaching and learning is incremental and ongoing, with innovations needing to be adapted to suit local needs [100, p.49, 67].
@@
“Philosophical inquiry as pedagogy must be integrated into the curriculum as well as integrate the curriculum.” [101, p.1057].
@@ 3.7 Implications of our Framework
“It is important that developers of quality frameworks not fall into the trap of substituting intuition and guesswork for evidence-based validation processes.” [111, p.361]
@@ 4 WoT Frameworks - A Comparative Study
“Each world whilst it is attended to is real after its own fashion; only the reality lapses with the attention.” [121, p.207].
@@ 4.2 English’s Framework
Systems Thinking and Design-Based Thinking are not discussed here to avoid redundancy. See Tables 2 and  4.
@@
(1) Critical Thinking: Several undergraduate STEM reform documents emphasize critical thinking as essential for developing and refining ideas, whether through explanations, designs, or investigations. Engaging in argumentation and critique foster this thinking and deepen understanding [44, p.125],  [43, p.xv], [79, p.46]. Given this emphasis in reform documents, critical thinking rightly finds a place in English’s framework.
@@
While research indicates that argumentation can enhance conceptual learning [124, p.1304, p.1305][125], Osborne expresses concern about its notable absence in science education. He highlights the need for a deeper understanding of how argumentation facilitates learning and the characteristics of learning environments that foster effective student arguments [57]. In our exploration of SWoT frameworks within scientific argumentation and evidence-based reasoning, we found Slavit et al.’s application of Toulmin’s Claim-Evidence-Reasoning (CER) framework [126, 127, 125] in the context of integrated STEM to be particularly insightful [24, p.135][57, p.464]. An examination of Slavit et al.’s analytic coding framework provides valuable insights with potential applications for future qualitative analyses, particularly in applying CER within integrated STEM contexts.
@@
English’s approach is echoed by Stillman et al., who stress the importance of learning environments that enhance students’ social-critical efficacy. They promote a socio-critical mathematical modeling cycle, allowing students to address problems relevant to their interests and communities while producing solutions that are socially significant [129, p.18]. This perspective aligns with current educational practices that emphasize real-world applications and social relevance in mathematics.
In our context, as an example, a catapult design project enables students to collaborate in launching a projectile to a specific range while applying principles of projectile motion. This activity incorporates mathematical modeling, allowing students to make assumptions about factors such as launch angles, launch speed, and drag force. They engage in critical discussions regarding material choices and design constraints, fostering argumentation as they justify their decisions. Before creating an actual model, students can run simulations in Python to measure projectile range, which facilitates reflection and iteration as needed. Furthermore, showcasing their models at a community event not only makes the project fun and educational but also extends the reach of science to the broader community.
@@
In this context, Gardner implores us to get comfortable with the truth, emphasizing that true inquiry transcends mere conversation. She cautions educators to differentiate between facilitating dialogue and facilitating genuine inquiry, noting that quite often “just dialogue continues to pass for philosophical inquiry” [131, p.72]. This distinction has prompted us to reflect on our practices critically.
@@ 5 Summary
To address RG2, we conducted an in-depth comparison between our framework and two others: one by Dalal et al. in the context of Engineering Education Research (EER), and the other by English, focusing on Mathematical and STEM-based problem solving. We outlined the context of each of the studies in sections (i) and (i), and elaborated on each framework in sections (ii) and (iii) respectively. It was insightful and illuminating to draw connections between our framework and the other two, particularly in the comparative analyses shown in Tables 4 and 5. While acknowledging the limitations of our comparative analysis in section 4.3, we also identified a few key implications of our comparative study in section 4.4.
@@ 6 Future Work
In the second installment of our two-part series, we will demonstrate the application of the WoT4EDP framework as an analytical lens to examine new data from student group discussions and written reports in student-generated engineering design projects. We will quote verbatim at least two examples of physics-based ED problems generated by students. This study would explore how students engage with design, science, mathematics, and metacognitive reflection throughout their problem-solving processes. Additionally, we investigate students’ demonstrations of computational thinking—particularly in their Python coding—as they tackle real-world challenges. We would like to add a caveat that the data was collected in Spring 2023, well before the lead author encountered the term SWoT frameworks (see Section 2.5).
A key challenge in qualitative research is the issue of overlapping codes, which we addressed extensively in our recently accepted manuscript. A common approach to qualitative analysis involves segmenting textual data, yet given the complexity of students’ reasoning and expression, multiple ways of thinking often co-occur within a single statement. While we previously illustrated such code co-occurrences with examples, qualitative analysis extends beyond mere segmentation. Over-segmentation, in particular, risks fragmenting students’ thought processes, raising critical questions about how best to delineate and interpret their ideas. We explored these issues in depth in our prior work and will expand on this discussion in the forthcoming study [21, p.7-10].
In the follow-up study, the complexity of the data increases as we analyze not only students’ written and verbal responses but also their Python code, images, graphs, and tables. Consequently, our analysis requires a more nuanced approach beyond textual segmentation. To address this, we employ a multi-pronged strategy integrating traditional coding [137], thematic analysis [47], and thick description [49, 48, 50]. We will revisit these challenges in the follow-up paper, providing explicit details on our data parsing methods and coding sheets.
While we present one approach to analyzing this complex dataset, we acknowledge that different researchers may adopt varied methodologies depending on their research goals and interests. By applying the WoT4EDP framework, we aim to generate insights into students’ Integrated Thinking and Learning Innovation, as emphasized by English. This study not only extends the application of our framework but also contributes to the broader discourse in STEM education on effective teaching and learning practices. In the process, we hope to contribute to the discussion on the use of qualitative methods in PER.